need to find the measure of an angle whose supplement measures 38 degrees less than three times its complement.
Answer by checkley77(7072)

180-X=3(90-X)-38
180-X=270-3X-38
-X+3X=270-180-38
2X=52
X=52/2
X=26 DEGREE IS THE ANGLE.
PROOF:
180-26=3(90-26)-36
154=3*64-38
154=192-38
154=154